[figure: 9.7 cm]

2D fermions near an antiferromagnetic instability: non Fermi liquid

behavior and spin-mediated pairing

Andrey V. Chubukov Department of Physics, University of Wisconsin, Madison, WI 53706

I plan to review recent works with Ar. Abanov, B. Altshuler, A. Finkelstein and J. Schmalian on the properties of 2D fermions near the quantum-critical point towards antiferromagnetism. This problem is relevant to both cuprates and heavy fermion materials. I show that the width of the Fermi liquid region progressively shrinks as the system moves towards quantum criticality, and there is a wide range of quantum-critical behavior. In the quantum-critical regime, the fermionic spectral function interpolates between marginal behavior at intermediate frequencies, and \sqrt{ω, T} behavior at high frequencies. I discuss the pairing problem in the quantum-critical regime and show that it is qualitatively different from BSC theory because the pairing involves fully incoherent fermions and retarded, nearly gapless pairing potential. I show that due to the interplay between fermionic incoherence and the softening of the pairing boson the pairing instability temperature T_{ins} saturates at a finite value when the magnetic correlation length diverges. I then show that immediately below T_{ins} fermions form singlet pairs, but still remain incoherent. The coherence is recovered only at a much smaller temperature T_{c} which, at least in some doping range, scales inversely with the magnetic correlation length. I argue that the region between T_{ins} and T_{c} is the pseudogap region. I discuss experimental signatures of these results for the cuprates.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : Andrey V. Chubukov

Address for correspondence : Department of Physics, University of Wisconsin, Madison, WI 53706

Email address : chubukov@skazka.physics.wisc.edu